OpenStudy (anonymous):

"A malted milk ball with a radius of (0.20 in) is coated with chocolate. What is the surface area of the chocolate? Ignore the thickness. Leave your answer in terms of π." The answer i got was 0.5024, is this correct?

OpenStudy (phi):

Leave your answer in terms of π. It looks like you multiplied by 3.14, which is not what they want.

OpenStudy (shane_b):

\[SA=4pir^2\]

OpenStudy (anonymous):

Not to be rude but I know the formula I just don't know if I did it right. Phi what do you mean when you say thats not what they want?

OpenStudy (phi):

the question says : leave your answer in terms of pi. that means they want 0.16 pi not 0.16*3.14= 0.5024

OpenStudy (anonymous):

So it would only be 0.16?

OpenStudy (phi):

not quite. the answer is \( 0.16 \pi \) remember that \( \pi \) is goes on forever, and 3.14 is only a close approximation. The exact answer leaves s the pi in the answer.

OpenStudy (phi):

Here are more digits of pi http://www.piday.org/million.php

OpenStudy (anonymous):

Ohh okay, I see what you mean. Thank you.